sxdsouls:

Yosef applies an input force of 50 N to a crowbar. The crowbar applies a force of 750 N to the lid of a crate. What is the mechanical advantage of the crowbar? The mechanical advantage of the crowbar is _.? idk its science

XxAngelGirlxX:

Explanation: 15 Given that, Input force applied by the Yosef to a crowbar,MA- \[\frac{ f 2}{ f1 }\] Output force applied by the crowbar, \[\frac{ 750 }{ 50 }\] We need to find the mechanical advantage of the crowbar. The mathematical formula of the mechanical advantage is given by : MA = output force/input force MA=\dfrac{F_2}{F_1} MA=\dfrac{750}{50} MA = 15 So, the mechanical advantage of the crowbar is 15. Hence, this is the required solution.